OpenStudy (anonymous):

how do I solve this for x...55.5= -0.3476x^2+10.948x-23.078

OpenStudy (aum):

subtract 55.5 from both sides. divide both sides by 0.3476. Solve the quadratic.

OpenStudy (anonymous):

that will be 0= -0.3476x^2+10.948x-78.578

OpenStudy (anonymous):

@Compassionate ...using the quadratic formular I'm getting an undefined answer.

OpenStudy (aum):

55.5 = 0.3476x^2 + 10.948x - 23.078 subtract 55.5 from both sides: 0 = 0.3476x^2 + 10.948x - 78.578 divide both sides by 0.3476: 0 = x^2 + 31.496x - 226.059 x^2 + 31.496x - 226.059 = 0 use quadratic formula to solve for x.

OpenStudy (aum):

\( \Large x = \frac{-31.496 \pm \sqrt{31.496^2 - 4(1)(-226.059)}}{2} = ? \)

OpenStudy (anonymous):

this is -0.3476x^2

OpenStudy (anonymous):

ok..let me calculate it

OpenStudy (aum):

The coefficient of x^2 in your original posting is +0.3476 not minus.

OpenStudy (anonymous):

yea..i made a mistake it is suppose to be minus

OpenStudy (aum):

You can always go back and edit the post.

OpenStudy (aum):

Then the equation will be: x^2 - 31.496x + 226.059 = 0 \( \Large x = \frac{31.496 \pm \sqrt{(-31.496)^2 - 4(1)(226.059)}}{2} = \\ \Large x = \frac{31.496 \pm 9.368}{2} = \frac{40.864}{2}, \frac{22.128}{2} = ? \)

OpenStudy (anonymous):

20.432,11.064

OpenStudy (aum):

Yes.

OpenStudy (aum):

Round the answers If they ask you to do so. Otherwise leave it up to 3 decimals because the original problem has coefficients up to 3 decimal places.